7.(9pts) Solve the initial value problem by the method of Laplace transform: y+ y = u(t - 3), y(O) = 0, y(0) = 1.
Answer #1

y000= yty ut-3) 빵 delee 매 II) yo) - 9.50) - le) + Yi) = 듯 (+)() -1 -10 지 로 + 9 = (81) ( py- []) 99 Sint + (1 - bect) dt)